Write extensively various methods use in determination of standard enthalpy
1
Expert's answer
2020-10-27T12:39:34-0400

Solution

Several methods are used to determine the standard enthalpy of a reaction. The most common of them is the measurement of the amount of heat that is released or absorbed in the calorimeter. To do this, use the formula:

"\Delta H = -\frac{Q}{n} = \frac{c(H2O) \times m(solution) \times \Delta T}{n}"

Another way to calculate the change in the enthalpy of a reaction is to use experimental values or theoretical values (in the table) of the enthalpy of formation of a substance using the following formula:

"\Delta H = \sum(\Delta H(prod.)) - \sum(\Delta H(reag))"

Another way to determine the change in the enthalpy of a reaction comes from determining the change in the enthalpy. The thermal effect of a chemical reaction depends only on the initial and final state of the system, and does not depend on the intermediate stages of the reaction. In this way, you can use other reactions involving those compounds that appear in the original (first) reaction equation.

Answer:

Several methods are used to determine the standard enthalpy of a reaction. The most common of them is the measurement of the amount of heat that is released or absorbed in the calorimeter. To do this, use the formula:

"\Delta H = -\frac{Q}{n} = \frac{c(H2O) \times m(solution) \times \Delta T}{n}"

Another way to calculate the change in the enthalpy of a reaction is to use experimental values or theoretical values (in the table) of the enthalpy of formation of a substance using the following formula:

"\Delta H = \sum(\Delta H(prod.)) - \sum(\Delta H(reag))"

Another way to determine the change in the enthalpy of a reaction comes from determining the change in the enthalpy. The thermal effect of a chemical reaction depends only on the initial and final state of the system, and does not depend on the intermediate stages of the reaction. In this way, you can use other reactions involving those compounds that appear in the original (first) reaction equation.